LBTB

02:41
  • Starting with ESM.
  • Bin2 400 Hz 600 EM gain
  • mag 14.2
  • g_tt_3.54_ho1_9.14_ho2_0.08.fits
03:18
  • New target.
  • Looking for one that is not into the wind.
  • Full AO
  • OK, the OCAM is still live after changing binning...
  • Something wrong with the script. Our pointing was good, but we didn't see anything.
03:29
  • New target
  • ESM
  • Bin 2 400 Hz
  • Mag 14.2
  • g_tt_5.38_ho1_9.29_ho2_0.08.fits
04:00
  • New target.
  • g_tt_0.66_ho1_0.34_ho2_0.10.fits
  • Bin 4 374 Hz
  • 15.5 mag
04:32 05:07
  • Re-closed after a big cloud.
  • bin4 390Hz
  • g_tt_0.74_ho1_0.73_ho2_0.10.fits
  • mag 15.6 half is moonlight in the cloud.
  • Closed due to wind.
06:50
  • Adsec Failure for no apparent reason. The dome is still closed.
  • Recovered without problem. We have still gusts of 22m/s
8:26
  • Re-opening
  • Going to ESM target.
  • Eventually......
  • "script issues"
  • Going.
  • 2022-05-12 08:33:29.107640| -13006 - [AOException] centerStar: WARNING - RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND
  • Looks like we are calling AO off. Cloudy and bad seeing.

-- GregoryTaylor - 12 May 2022
Topic revision: r3 - 12 May 2022, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback